# Gaussian process modelling of infectious diseases using the Greta software package and GPUs

### v1

@@ Abstract
Gaussian process are a widely-used statistical tool for conducting non-parametric inference in applied sciences, with many computational packages available to fit to data and predict future observations. We study the use of the Greta software for Bayesian inference to apply Gaussian process regression to spatio-temporal data of infectious disease outbreaks and predict . Greta builds on Tensorflow, making it comparatively easy to take advantage of the significant gain in speed offered by GPUs. In these complex spatio-temporal models, we show a reduction of up to 70% in computational time relative to fitting the same models on CPUs. We show how the choice of covariance kernel impacts the ability to infer spread and extrapolate to unobserved spatial and temporal units. The inference pipeline is applied to weekly incidence data on tuberculosis in the East and West Midlands regions of England over a period of two years.
@@ 1 Introduction
Whilst their analytical tractability facilitates fitting of GPs, inference and prediction requires complex linear algebra operations of often high-dimensional matrices, making computation slow and cumbersome. A wide variety of approaches have been developed to improve efficiency and enable practitioners to approach larger scale problems common to infectious disease modelling.
Previous authors have shown the wide variety of approaches that GPs can be used for in infectious disease modelling. For example, keelinggps (1) used GPs to estimate growth rates, $r$, across an area of England in the SARS-CoV-2 pandemic, using a Matèrn covariance kernel. The flexible
Other authors use GPs to emulate the behaviour of mechanistic models, such as SIR models and their variants. Based on careful runs at training points in the parameter space, the GP emulators can interpolate the behaviour of the mechanistic model at other to provide a computationally cheap approximation to the model surface. DUNNE2022 (2) use GP emulators to calibrate the mean and variance of an age-structured compartmental model of asymptomatic spread of SARS-CoV-2 in hospitals, whilst BUCKINGHAMJEFFERY2018111 (3) and trostle24 (4) directly model a moment-closure approximation to the stochastic jump process of an SIR model, with the latter extending to a spatial context.
In section 2, we introduce the GP model structure for machine learning and the Bayesian workflow required for successful fitting, assessment and prediction using the R package greta.gp utilising Tensorflow on a GPU. In 3, we apply the methodological workflow to a case study of Tuberculosis incidence in the Midlands region of England. Finally in Section 4, the implications of the work and areas for further extensions are discussed.
@@ 2 Gaussian processes
A Gaussian Process is a collection of random variables, any finite number of which have (consistent) joint Gaussian distributions (5). Gaussian processes are often used to model a smooth function $f$ of data $X$. In this case, the collection of random variables is the set $\{f(x):x\in X\}$, and is a Gaussian Process provided any finite subset, $\{f(x_{1}),\ldots,f(x_{n})\}$ for some $x_{1},...,x_{n}\in X$, has a multivariate normal distribution with mean function $m(\cdot)$ and kernel function $k(\cdot,\cdot)$. That is:
@@ 4. Periodic kernel
A useful property of kernel functions is that both the sum and product of two kernels is also a kernel (5). This makes modelling spatio-temporal processes easier as, assuming separability, we can assign different kernel functions to the space and time components of the model. Additive kernel functions are generally used for modelling the independent effects of the explanatory variables, whilst multiplicative kernels are generally used for the interactive effect between variables.
@@ 2.2 Combined kernel functions
This type of approach has been used in previous research, as by (e.g., 6, 7, 8).
@@ 2.3.1 Inferring parameters
The parameters of the response distribution, as well as the GP mean and kernel functions need to be estimated from the data. The most common ways are maximum likelihood estimation (e.g., 9) and Bayesian inference (e.g., 10). Using a Bayesian Markov chain framework is more computationally intensive because of the necessity to repeatedly invert the covariance matrix to obtain the posterior distribution, however it provides a straightforward means of assessing parameter uncertainty through the posterior distribution. Understanding uncertainty is particularly important in an epidemic context to enable associated risks to be taken into consideration. This uncertainty naturally reduces in areas where data are particularly informative, whilst allowing forecasts to be obtained with associated uncertainties.
@@ 2.3.2 Model checking and comparison
In order to assess fitted models alignment to the data and within and out of sample prediction, posterior predictive checks can be conducted. These approaches compare the density of the predictive and observed values, and secondly by applying a discrepancy measure on observed and simulated data to see if they give similar values. In the latter case we use the Freeman-Tukey statistic (2.3.2), which is less sensitive to smaller values than for example the Chi-Squared statistic (11). If the model is a good fit then a plot of the observed and simulated Freeman-Tukey values they should show even spread over the $y=x$ line. The Freeman-Tukey statistic is calculated as:
$T(\boldsymbol{y},\boldsymbol{\theta})=\sum_{i}(\sqrt{y_{i}}-\sqrt{E(y_{i}|\boldsymbol{\theta}}))^{2}$
To compare the models, three metrics designed for Bayesian inference will be utilised: leave-out-one cross validation, continuous ranked probability score and a Bayesian p-value. They each serve different purposes in model comparison. Our first metric was an approximate leave-out-one cross validation (LOO-CV). LOO-CV conducted over all posterior samples and data points would be very time consuming, so approximations can be made, which use the log-likelihood of the data given posterior draws of the distribution’s parameters. Such an approximation is used by the loo package in R, and gives an information criteria ‘looic’. As with other information criteria, the lower the looic score the better the model fits the data. looic records how well the model performs when making predictions on the training data.
The second metric, continuous ranked probability score (CRPS), considers how well the model can predict new data. If the model is a good fit, one would expect the CDF of the predictions to be similar to a function that is 1 if the prediction is greater than or equal to the observed value and 0 otherwise (12). If the model is good at making predictions then the difference in area between these two functions should be close to 0.
@@
This was shown by (13) to be equivalent under certain conditions (finite first moment) to:
@@
The third metric is the Bayesian P-value. It is a formalisation of the posterior predictive check using the discrepancy measure. From (14) the Bayesian p-value is defined as:
@@
where T is the test statistic and the probability is taken over the posterior distribution of $\boldsymbol{\theta}$ and the posterior predictive distribution of $y$ ($y_{rep}$). A good p-value is one that is close to 0.5; although anywhere between 0.05 and 0.95 is deemed acceptable (14). Here we again use the Freeman-Tukey test statistic as $T$.
@@ 2.4 Prediction
Gaussian Processes make generating predictions relatively simple since they have a tractable form for the posterior predictive distribution. If the observed data is the matrix $\mathbf{X}$ and the predicted new data is $\mathbf{X_{*}}$, such that $\mathbf{f}$ and $\mathbf{f}_{*}$ is the function on the old and new points respectively, then the joint distribution is (5):
@@
$\mathbf{f}_{*}|\mathbf{f},\mathbf{X},\mathbf{X}_{*}\sim N(m(\mathbf{X}_{*})+\mathbf{K}_{*}^{T}\mathbf{K}^{-1}(\mathbf{f}-m(\mathbf{X}),\quad\mathbf{K}_{**}-\mathbf{K}_{*}^{T}\mathbf{K}^{-1}\mathbf{K}_{*}).$
@@ 2.5 Approximations
The use of Gaussian Processes in practice can be limited by the computational complexity of calculating the inverse for the kernel, which increases with $O(n^{3})$ as $n$ (the number of training cases) increases (15). The inversion is necessary to make predictions and calibrate parameters. Hence, when the data is large (more than $\sim$ 2500) (6) it is often best to take a subset of $m<n$ data points, which is known as a sparse approximation. How this subset is used depends on the approximation method.
There are several methods of sparse approximation. The baseline approximation is the Subset of Data (SoD) Approximation. This method simply takes a subset $m$ of $n$ of the training data. This reduces the computational complexity to $O(m^{3})$, but ignores a lot of the data. The method that greta.gp has, and that used in this analysis, is the Subset of Regressors (SoR) Approximation. ubset of regressors was proposed by (16, 17) and adapted by (18). It is a reduced rank method, that is it approximates the original kernel functions by a weighted sum of the kernel functions applied to the subset. Let the original n functions be written $S_{\mathcal{N}}=\{k(x,x_{i}|\theta):i\in\mathcal{N}\}$ with $\mathcal{N}=\{1,2,...,n\}$. If we select the inducing variables (indices) $\mathcal{M}\subset\mathcal{N}$ such that $S_{\mathcal{M}}=\{k(x,x_{j}|\theta):j\in\mathcal{M}\}\subset S_{\mathcal{N}}$, then we write the approximation as $\hat{k}(x,x_{r}|\theta)=\sum_{j\in\mathcal{M}}c_{jr}k(x,x_{j})$ for $r\in\mathcal{N}$. The coefficients are found by minimising the error function. A disadvantage of this approximation, is that it may underestimate the predictive variance when a data point is far away from points in the subset (15). See (15) for details on other common approximations.
The subset of the inputs used (inducing inputs) need to be chosen carefully as these change the final solution (while the inducing variables do not) (15). Three common ways of choosing inducing inputs are greedy selection, grid selection, and K-means ++ (19). Greedy selection, suggested by (20), adds points iteratively if they give the best fit in a batch. This method is extremely computationally expensive, as it involves fitting the Gaussian Process many times. K-Means ++ is good at taking into account the structure of the data, but is also computationally intensive.
@@ 2.5.1 Use of Greta software in R
The GPs are fitted in R (21) using the package greta (22), which was designed with three main improvements in mind over existing software (23): analysis can be carried out directly in R; easy to extend the package in R; and fast inference for large datasets using TensorFlow (24). In this paper, we also used an extension of the greta package, greta.gp (25), which facilitates the fitting of GPs.
There are currently three sampling algorithms available in greta: Random walk Metropolis-Hastings, Hamiltonian Monte Carlo, and a multivariate slice sampling algorithm. Random Walk can struggle with high dimension probability distributions, which is inherent in Gaussian Processes (26). Although, slice sampling has been used for Gaussian Processes previously by (27), we found that it sampled poorly when we trialled it. So we chose to use Hamiltonian Monte Carlo for all our models. This method requires you to select a minimum and maximum number of leapfrog steps. If the number of leapfrogs is low, samples will have high autocorrelation, but if high, it takes much longer to sample (28). We found for our models setting the minimum number to 15 and the maximum number 20 was a good trade-off between a slightly slower model and sampling well.

### v3

@@ Abstract
Gaussian process are a widely-used statistical tool for conducting non-parametric inference in applied sciences, with many computational packages available to fit to data and predict future observations. We study the use of the Greta software for Bayesian inference to apply Gaussian process regression to spatio-temporal data of infectious disease outbreaks and predict future spread. Greta builds on Tensorflow, making it comparatively easy to take advantage of the significant gain in speed offered by GPUs. In these complex spatio-temporal models, we show a reduction of up to 70% in computational time relative to fitting the same models on CPUs. We show how the choice of covariance kernel impacts the ability to infer spread and extrapolate to unobserved spatial and temporal units. The inference pipeline is applied to weekly incidence data on tuberculosis in the East and West Midlands regions of England over a period of two years.
@@ 1 Introduction
Whilst their analytical tractability facilitates fitting of GPs, inference and prediction requires complex linear algebra operations of often high-dimensional covariance matrices, making computation slow and cumbersome for many realistic data. A wide variety of approaches have been developed to improve efficiency and enable practitioners to approach larger scale problems common to infectious disease modelling (e.g., Ambikasaran et al., 2016; Moran and Wheeler, 2022, and references therein). Approaches to improve computation reduce the overhead of the problem by providing approximately independent subsets of the data through sparse approximations to the covariance kernel. The linear algebra operations can then be applied repeatedly on smaller subsets, a task that is well-suited to both parallelisation and fast computation on Graphical Processing Units (GPUs). Careful selection of these subsets can reduce unwarranted approximation errors in this.
Previous authors have shown the wide variety of approaches that GPs can be used for in infectious disease modelling. For example, Guzmán-Rincón et al. (2023) used GPs to estimate growth rates, $r$, across an area of England in the SARS-CoV-2 pandemic, using a Matèrn covariance kernel. Reed et al. (2024) embed a GP as a latent process within a Poisson regression to interpolate snail distributions and allowing risks of schistomiasis to be estimated in Malawi.
Other authors use GPs to emulate the behaviour of mechanistic models, such as SIR models and their variants. Based on careful runs at training points in the parameter space, the GP emulators can interpolate the behaviour of the mechanistic model at other to provide a computationally cheap approximation to the model surface. Dunne et al. (2022) use GP emulators to calibrate the mean and variance of an age-structured compartmental model of asymptomatic spread of SARS-CoV-2 in hospitals, whilst grapKINGHAMJEFFERY2018111 and Trostle et al. (2024) directly model a moment-closure approximation to the stochastic jump process of an SIR model, with the latter extending to a spatial context. Relatively little work has been done previously on the use of GPUs for infectious disease models, both due to the cost of the hardware making them prohibitively expensive to the average researcher, and also a steep learning curve in making general purpose software difficult to implement on them. There is also not a guarantee that they will provide significant improvement in all scenariosFagard-Jenkin and Thomas (2024). However, costs have reduced and they are now more readily available, with new software allowing integration of standard algorithms. Where these hardware have been used in infectious disease models, it has largely been in mechanistic individual based models where the GPU simulations enable repetition of computations across individuals giving significant speed improvements(e.g., Galvão Filho et al., 2016; Leonenko et al., 2015).
In section 2, we introduce the GP model structure for machine learning and the Bayesian workflow required for successful fitting, assessment and prediction using the R package greta.gp utilising tensorflow-probability on a GPU. In 3, we apply the methodological workflow to a case study of Tuberculosis incidence in the Midlands region of England. Finally in Section 4, the implications of the work and areas for further extensions are discussed.
@@ 2 Gaussian processes
A Gaussian Process is a collection of random variables, any finite number of which have (consistent) joint Gaussian distributions Rasmussen (2004). Gaussian processes are often used to model a smooth function $f$ of data $X$. In this case, the collection of random variables is the set $\{f(x):x\in X\}$, and is a Gaussian Process provided any finite subset, $\{f(x_{1}),\ldots,f(x_{n})\}$ for some $x_{1},...,x_{n}\in X$, has a multivariate normal distribution with mean function $m(\cdot)$ and kernel function $k(\cdot,\cdot)$. That is:
@@ 4. Periodic kernel
Figure 1: Comparison of smoothness of varying kernels (Matèrn, radial basis function and periodic) for some synthetic data.
@@
A useful property of kernel functions is that both the sum and product of two kernels is also a kernel (Rasmussen, 2004). This makes modelling spatio-temporal processes easier as, assuming separability, we can assign different kernel functions to the space and time components of the model. Additive kernel functions are generally used for modelling the independent effects of the explanatory variables, whilst multiplicative kernels are generally used for the interactive effect between variables.
@@ 2.2 Combined kernel functions
This type of approach has been used in previous research, as by (e.g., Senanayake et al., 2016; Albinati et al., 2017; Hawryluk et al., 2021).
2.3 Observation process
In many situations, the Gaussian process is capable of capturing the underlying spatio-temporal latent structure in the data, however is inappropriate to directly model the observed data. For example, disease incidence is often measured through counts of new infections, a discrete random variable. Therefore, it is necessary to embed the Gaussian process within a probability mass (discrete) or density (continuous) function to map from a continuous valued output to an appropriate support, as well as account for additional errors or uncertainties in the data collection.
In larger populations, we would generally expect a higher absolute number of infections. As a result, it is common to split the mean number of infections ($\mu_{ij}$) as a product of the background population effect ($e_{ij}$) and the excess risk ($\theta_{ij}$) (Lawson, 2021) as to focus on modelling the excess risk:
$\mu_{ij}=e_{ij}\theta_{ij}.$ (2)
$e_{ij}=p_{ij}R$
$R=\frac{\sum_{i}\sum_{j}y_{ij}}{\sum_{i}\sum_{j}p_{ij}}$
$\log(\theta_{ij})=f(x_{ij})$
$f(.)\sim GP(m(.),k(.,.)).$
The relationship between the data and the Gaussian process is hence:
$y_{ij}\sim\text{Pdf}(\mu_{ij},\phi_{ij}),$ (3)
where $Pdf(.)$ is the selected observation response distribution with mean $\mu_{ij}$, as defined as in equation 2, and $\phi_{ij}$ are additional distribution-specific parameters, such as variance or dispersion parameters.
@@ 2.4.1 Inferring parameters
Once the model structure in equation (3) has been defined, the parameters of the response distribution ($\phi_{i,j}$), as well as the GP mean and kernel functions need to be estimated from the data. The most common ways are maximum likelihood estimation (e.g., Mardia and Marshall, 1984) and Bayesian inference (e.g., MacKay, 1992), the latter requiring additional specification of prior distributions for the model parameters. Using a Bayesian Markov chain Monte Carlo framework (see Andrieu et al., 2003, for more details) is more computationally intensive because of the necessity to repeatedly invert the covariance matrix to obtain the posterior distribution, however it provides a straightforward means of assessing parameter uncertainty through the posterior distribution. Understanding uncertainty is particularly important in an epidemic context to enable associated risks to be taken into consideration. This uncertainty naturally reduces in areas where data are particularly informative, whilst allowing forecasts to be less informed in areas of space and/or time of data sparsity.
@@ 2.4.2 Model checking and comparison
In order to assess fitted models alignment to the data and within and out of sample prediction, posterior predictive checks can be conducted. These approaches compare the density of the predictive and observed values, and secondly by applying a discrepancy measure on observed and model-fitted data or observed and data simulated from the fitted model to see if they give similar values. In the latter case we use the Freeman-Tukey statistic (2.4.2), which is less sensitive to smaller values than for example the Chi-Squared statistic (Conn et al., 2018). If the model is a good fit then a plot of the observed and simulated Freeman-Tukey values should show even spread over the $y=x$ line. The Freeman-Tukey statistic is calculated as:
$T(\boldsymbol{y},\boldsymbol{\theta})=\sum_{i}(\sqrt{y_{i}}-\sqrt{E(y_{i}|\boldsymbol{\theta}}))^{2},$
where $y_{i}$ are the observed data and $E(y_{i}|\boldsymbol{\theta})$ are the fitted model values for those data.
To compare between models with differing response distributions or kernel functions, three metrics designed for Bayesian inference will be utilised: leave-out-one cross validation, continuous ranked probability score and a Bayesian p-value. They each serve different purposes in model comparison. Our first metric was an approximate leave-out-one cross validation (LOO-CV) Vehtari et al. (2017). This approach aims to fit the model to a subset of the data and use this fitted model to predict the removed observations. LOO-CV conducted over all posterior samples and data points would be very time consuming, so approximations can be made, which use the log-likelihood of the data given posterior draws of the distribution’s parameters. Such an approximation is used by the loo package in R, and gives an information criteria ‘looic’ and is described in Algorithm 1. As with other information criteria, the lower the looic score the better the model fits the data. looic records how well the model performs when making predictions on the training data.
Algorithm 1 Loo Algorithm
1:Make S = 200 posterior simulations of f, $\phi$ and $\lambda$
2:Make an empty matrix of dimension S by N (# data points = 6760)
3:for $i$ in 1:S do
4:     for $j$ in 1:N do
5:         Calculate $\mu$ using equation 2
6:         Convert $\mu$ and $\phi$ to $r$ and $p$ using equations 3.2 and 3.2
7:         Calculate $\pi$ using equation 3.3
8:         Calculate log-likelihood of observed values given $\pi$, $r$ and $p$
9:         Store to position i, j in the matrix
10:Calculate relative effective sample sizes from the log-likelihood matrix using relative_eff from loo package.
11:Calculate loo using the relative effective sample sizes and log-likelihood matrix using loo from loo package.
The second metric, continuous ranked probability score (CRPS), considers how well the model can predict new data. If the model is a good fit, one would expect the CDF of the predictions to be similar to a function that is 1 if the prediction is greater than or equal to the observed value and 0 otherwise Faran (2023). If the model is good at making predictions then the difference in area between these two functions should be close to 0.
@@
This was shown by Gneiting and Raftery (2007) to be equivalent under certain conditions (finite first moment) to:
@@
Algorithm 2 CRPS Algorithm
1:Make 1000 predictions for each location for each week 105 to 108
2:for week in 105:108 do
3:     $y\_pred\leftarrow$ predictions[,time ==absent$=$ = week]
4:     $X\leftarrow$ $y\_pred$[1:500,] $\triangleright$ split into two predictive samples
5:     $X^{\prime}\leftarrow$ $y\_pred$[,501:1000]
6:     Calculate CRPS using X, X’ and y (the observations) via the crps function from the loo package.
The third metric is the Bayesian P-value, Algorithm 3, aims to determine whether predictive data simulated from the fitted model shows similar characteristics to the osberved data. It is a formalisation of the posterior predictive check using a specified discrepancy measure. From Gelman et al. (2013) the Bayesian p-value is defined as:
@@
where T is the test statistic and the probability is taken over the posterior distribution of $\boldsymbol{\theta}$ and the posterior predictive distribution of $y$ ($y_{rep}$). A good p-value is one that is close to 0.5; although anywhere between 0.05 and 0.95 is deemed acceptable (Gelman et al., 2013). Here we again use the Freeman-Tukey test statistic as $T$.
Algorithm 3 Bayesian P-value
1:Tukey.obs $\leftarrow$ Freeman-Tukey Statistic (equation 2.4.2) in terms of the observed and expected value
2:Tukey.sim $\leftarrow$ Freeman-Tukey Statistic in terms of the simulated response and expected value
3:Make 1000 simulations of Tukey.obs and Tukey.sim using posterior parameter draws
4:Bayes_p $\leftarrow$ mean(Tukey.obs$>$Tukey.sim)
@@ 2.5 Prediction
Gaussian Processes make generating predictions relatively simple since they have a tractable form for the posterior predictive distribution. If the observed data is the matrix $\mathbf{X}$ and the predicted new data is $\mathbf{X_{*}}$, such that $\mathbf{f}$ and $\mathbf{f}_{*}$ is the function on the old and new points respectively, then the joint distribution is (Rasmussen, 2004):
@@
$\mathbf{f}_{*}|\mathbf{f},\mathbf{X},\mathbf{X}_{*}\sim N(m(\mathbf{X}_{*})+\mathbf{K}_{*}^{T}\mathbf{K}^{-1}(\mathbf{f}-m(\mathbf{X})),\quad\mathbf{K}_{**}-\mathbf{K}_{*}^{T}\mathbf{K}^{-1}\mathbf{K}_{*}).$
@@ 2.6 Approximations
The use of Gaussian Processes in practice can be limited by the computational complexity of calculating the inverse for the kernel, which increases with $O(n^{3})$ as $n$ (the number of training cases) increases (Quiñonero-Candela and Rasmussen, 2005). The inversion is necessary to make predictions and calibrate parameters. Hence, when the data is large (more than $\sim$ 2500) (Senanayake et al., 2016) it is often best to take a subset of $m<n$ data points, which is known as a sparse approximation. How this subset is used depends on the approximation method.
There are several methods of sparse approximation. The baseline approximation is the Subset of Data (SoD) Approximation. This method simply takes a subset $m$ of $n$ of the training data. This reduces the computational complexity to $O(m^{3})$, but ignores a lot of the data. The method that greta.gp has, and that used in this analysis, is the Subset of Regressors (SoR) Approximation. Subset of regressors was proposed by Wahba (1990); Poggio and Girosi (1990) and adapted by Smola and Bartlett (2000). It is a reduced rank method, that is it approximates the original kernel functions by a weighted sum of the kernel functions applied to the subset. Let the original n functions be written $S_{\mathcal{N}}=\{k(x,x_{i}|\theta):i\in\mathcal{N}\}$ with $\mathcal{N}=\{1,2,...,n\}$. If we select the inducing variables (indices) $\mathcal{M}\subset\mathcal{N}$ such that $S_{\mathcal{M}}=\{k(x,x_{j}|\theta):j\in\mathcal{M}\}\subset S_{\mathcal{N}}$, then we write the approximation as $\hat{k}(x,x_{r}|\theta)=\sum_{j\in\mathcal{M}}c_{jr}k(x,x_{j})$ for $r\in\mathcal{N}$. The coefficients are found by solving the equation $k(x_{j},x_{j})^{1/2\intercal}\mathbf{c}=S_{\mathcal{M}}$. A disadvantage of this approximation, is that it may underestimate the predictive variance when a data point is far away from points in the subset (Quiñonero-Candela and Rasmussen, 2005). See (Quiñonero-Candela and Rasmussen, 2005) for details on other common approximations.
The subset of the inputs used (inducing inputs) need to be chosen carefully as these change the final solution (while the inducing variables do not) (Quiñonero-Candela and Rasmussen, 2005). Three common ways of choosing inducing inputs are greedy selection, grid selection, and K-means ++ (Galy-Fajou and Opper, 2021). Greedy selection, suggested by Titsias (2009), adds points iteratively if they give the best fit in a batch. This method is extremely computationally expensive, as it involves fitting the Gaussian Process many times. K-Means ++ is good at taking into account the structure of the data, but is also computationally intensive.
@@ 2.6.1 Use of Greta software in R
The GPs are fitted in R (R Core Team, 2023) using the package greta (Golding, 2019a), which was designed with three main improvements in mind over existing software (Golding, 2019b): analysis can be carried out directly in R; easy to extend the package in R; and fast inference for large datasets using TensorFlow (Abadi et al., 2015). In this paper, we also used an extension of the greta package, greta.gp (Golding, 2024), which facilitates the fitting of GPs. Additions to the package were required, to both enable the zero-inflated probability mass functions appropriate for the data, as well as extensions to some of the kernel functions to provide the required flexibility. The periodic kernel did not allow you to choose which covariate to act on as it had only previously been used on time series data. The Tensorflow code was extended the code so that it can be applied to spatio-temporal data. These changes can be accessed through the forked versions of greta and greta.gp on the author’s GitHub. See the Supporting Information for instructions on how to make your own changes to greta and greta.gp.
There are currently three sampling algorithms available in greta: Random walk Metropolis-Hastings, Hamiltonian Monte Carlo, and a multivariate slice sampling algorithm. Random Walk can struggle with high dimension probability distributions, which is inherent in Gaussian Processes (Titsias et al., 2011). Although, slice sampling has been used for Gaussian Processes previously by Murray and Adams (2010), we found that it sampled poorly when we trialled it. So we chose to use Hamiltonian Monte Carlo for all our models. This method requires you to select a minimum and maximum number of leapfrog steps. If the number of leapfrogs is low, samples will have high autocorrelation, but if high, it takes much longer to sample (Arakawa et al., 2019). We found for our models setting the minimum number to 15 and the maximum number 20 was a good trade-off between a slightly slower model and sampling well.